OpenStudy (swissgirl):

Prove by induction: For all \[ n \in N, n+3 < 5n^2 \]

OpenStudy (swissgirl):

Mt textbook provided steps but i didnt follow

OpenStudy (swissgirl):

the first step was that it showed that this was true for 1

OpenStudy (amistre64):

n + 3 < 5n^2 id prolly rewrite it as 0 < 5n^2 - n -3

OpenStudy (amistre64):

induction ... never liked it ... says provide a basis step

OpenStudy (amistre64):

its true for some value; then assume its true for some general value "k"

OpenStudy (amistre64):

0 < 5k^2 - k -3 now prove its true for (k+1) 0 < 5(k+1)^2 - (k+1) -3

OpenStudy (swissgirl):

sorry my dad just pulled me away from the screen for a sec here

OpenStudy (swissgirl):

ill tell u the steps they gave

OpenStudy (swissgirl):

so first they showed its true for teh lowest vlaue of n=1 then they assumed that \( n \in N, n+3<5n^2 \)

OpenStudy (amistre64):

0 < 5(k+1)^2 - (k+1) -3 0 < 5(k^2 +2k +1) - k -1 - 3 0 < 5k^2 +10k +5 - k -1 - 3 rearrange 0 < (5k^2 -k -3) + (10k +5 -1) ^^^^^^ ^^^ this part is always positive this part is assumed true already

OpenStudy (swissgirl):

Then (n+1)+3=n+3+1 < \( 5n^2+1\) < \( 5n^2+10n+5 \) < \( 5(n+1)^2 \)

OpenStudy (swissgirl):

Ur proof is correct but i am suppossed to show that it is true for K+1 or n+1 whtvr

OpenStudy (amistre64):

right, after assuming its correct for n=k; then you readdress if for n=(k+1)

OpenStudy (swissgirl):

Like id ont get what was done by the proof i provided

OpenStudy (amistre64):

as long as you get your original "kth" value back and a value that cant possibly contradict the setup your good

OpenStudy (swissgirl):

ya but can u explain me the proof i gave. like i didnt follow it

OpenStudy (amistre64):

you said first step (called a basis) is to determine its value for a specific integer: 1 in this case n+3 < 5n^2 1 + 3 < 5 , is true

OpenStudy (swissgirl):

right that i followed

OpenStudy (amistre64):

second step is to assume its true for all n=k values k+3 < 5k^2

OpenStudy (swissgirl):

right

OpenStudy (amistre64):

since its true for "k" we want to prove its also true for "k+1" so that it follows out indefinantly

OpenStudy (swissgirl):

right

OpenStudy (amistre64):

(k+1) +3 < 5(k+1)^2 k+1 +3 < 5k^2 +10k + 5 we need to rewrite this so that we have our original "k" setup

OpenStudy (swissgirl):

ok

OpenStudy (amistre64):

(k+3) + 1 < (5k^2) + 10k + 5 - 1 - 1 ------------------------------ (k+3) < (5k^2) + 10k + 4

OpenStudy (amistre64):

since this is an inequality, we would have to see if 10k+4 is always postive; since increasing 5k^2 only makes it bigger

OpenStudy (swissgirl):

well it is since its all natural number

OpenStudy (amistre64):

correct, then depending on how much your teacher wants you to reinvent the wheel, this is proof enough

OpenStudy (swissgirl):

But like y did the textbook prove like the way i showed u. Then (n+1)+3=n+3+1 < 5n^ 2 +1 < 5n^ 2 +10n+5 < 5(n+1)^ 2

OpenStudy (amistre64):

or you could go further and say 10k+4 > 0 -4 -4 ----------- 10k > -4 /10 /10 ------------ k > -2/5 and since k is always greater than -2/5 .... its good

OpenStudy (amistre64):

im not sure why the author of your textbook would choose one way over another

OpenStudy (swissgirl):

whtvr ok

OpenStudy (swissgirl):

THHANKKKSSS

OpenStudy (amistre64):

lol, youre welcome :)